Use mathematical induction to prove that:
1+2+3+...+n=[n(n+1)]/2

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

Use mathematical induction to prove that:
1+2+3+...+n=[n(n+1)]/2
-----
Show it is true for n = 1::
1 = [1(1+1)]/2 = 2/2 = 1
----------------------
Assume it is true for n = k::
1+2+3+...+k = [k(k+1)]/2
----
Prove it is true for n = k+1::
1 + 2 + 3 +...+ k + (k+1) = [k(k+1)]/2 + (k+1)
-------
= [(k^2+k]/2 + (k+1)
------
= (k^2 + k + 2k + 2)/2
------
= (k^2 + 3k + 2)/2
------
= [(k+1)(k+2)]/2
------
= [(k+1)((k+1)+1)]/2
-------
QED
-------------
Cheers,
Stan H.
---------------

Answer by ikleyn(52781)

.
See the lesson
    - Mathematical induction and arithmetic progressions
in this site.

Also, you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic
"Method of Mathematical induction".